# Higgs Phases and Boundary Criticality in memoriam Peter Higgs

### v1

@@ IV.2.1 Open Boundary Duality Transformation
Next, consider the $\cramped{\kappa\to 0}$ ($\cramped{\kappa^{\prime}\to\infty}$) limit. Electric charges are turned off, restoring the electric 1-form symmetry and reducing to a pure U(1) gauge theory. In the dual theory the gauge field $\tilde{H}$ is turned off, $\cramped{\mathrm{d}\tilde{H}=0}$. The theory reduces to a Coulomb gas of magnetic monopole worldlines [26, 58], which has a transition separating the deconfined phase (low temperature condensate) and the confined phase (high temperature gas).
Lastly, consider the $\cramped{\beta\to 0}$ ($\cramped{\beta^{\prime}\to\infty}$) limit, the strong coupling limit of the original theory. In the electric theory, we can fix unitary gauge to remove $\theta$ and obtain the action $-\kappa\cos(A)$ on every link independently, so the system is fully disordered. Equivalently, if we use Eq. 103, setting $\beta=0$ forces all of the Bessel functions to vanish except when $e=j_{\mathrm{e}}=0$, which reduces the partition function to $\prod_{\ell}I_{0}(\kappa)$. In the dual theory, $\beta^{\prime}\to\infty$, the weak coupling limit, and we have the constraint $\cramped{\tilde{H}=-\mathrm{d}\tilde{A}_{\mathrm{m}}}$, which further implies $\cramped{\mathrm{d}\tilde{H}=0}$. The resulting theory then just has Lagrange multipliers that force the charge loops and membranes to vanish, so the theory trivializes completely.
IV.2.2 Dual Boundary Symmetry Breaking
We now consider setting $\kappa^{\prime}=0$ in Eq. 108 ($\kappa\to\infty$). The variables in play are $\tilde{A}_{\mathrm{m}}$ and $\tilde{j}_{\mathrm{m}}$, defined on all links of the dual lattice, and $\tilde{H}$ and $\tilde{b}$ on all plaquettes. In the bulk, the $\kappa^{\prime}=0$ means that the kinetic term for $\tilde{H}$ drops out and so the fluxes of $\tilde{H}$ (corresponding to electric charges of the original action) are completely unconstrained. In other words, the bulk $\tilde{H}$ is in the strong coupling limit. We may integrate out $\tilde{H}$, to obtain
@@ IV.2.2 Dual Boundary Symmetry Breaking
Now consider an open boundary. Recall that the boundary of the dual lattice is “flat”, as shown in Fig. 10, i.e. it has no cells extending into the vacuum. This means that no magnetic charge or magnetic flux can exit the system. More concretely, the constraint $\cramped{j_{\mathrm{e}}|_{\partial X}=0}$, in Eq. 102, in the electric variables is reflected in the dual constraint $\cramped{\mathrm{d}\tilde{H}|_{\partial\tilde{X}}=0}$ on every dual boundary link, which enforces that $\tilde{H}$ is pure gauge in the dual boundary layer. This means that the boundary action has no $\kappa$ dependence. We resolve the boundary constraint as $\tilde{H}=\mathrm{d}\tilde{f}$, so that
@@
We then define a composite field $\tilde{\chi}=\tilde{A}_{\mathrm{m}}+\tilde{f}$, which is gauge invariant under the gauge transformations $\tilde{A}_{\mathrm{m}}\to\tilde{A}_{\mathrm{m}}+\lambda$ and $f\to f-\lambda$. This allows us to rewrite the boundary action as
@@
To proceed from here we need to be careful about how the $\tilde{j}_{\mathrm{m}}$ and $\tilde{b}$ can move between the bulk and boundary layers. We do so in the $\kappa\to\infty$ limit, by combining this with the bulk action Eq. 111. We can then integrate out $\tilde{A}_{\mathrm{m}}$ to generate the magnetic Gauss law, which is enforced on every link. This leaves us with the total action
$S_{\text{dual}}\xrightarrow{\kappa\to\infty}-\frac{1}{2\beta^{\prime}}\langle\tilde{b},\tilde{b}\rangle_{\tilde{X}}-\frac{\beta^{\prime}}{2}(\mathrm{d}\tilde{\chi})_{\partial\tilde{X}}^{2}-i\langle\tilde{\chi},j_{\mathrm{m}}|_{\partial\tilde{X}}\rangle.$ (115)
@@
One should be concerned here as to how the $\tilde{j}_{\mathrm{m}}$ from the bulk (edges of $\tilde{b}$) couple to the boundary. The key to understand what happens here is that (i) the $\tilde{b}$ membranes can lie in the boundary layer where they cost zero action, and (ii) this action to be gauge-invariant under shifts of $\tilde{\chi}$, we must have an additional boundary Gauss law,
$\mathrm{d}^{\dagger}j_{\mathrm{m}}|_{\partial\tilde{X}}=0$ (116)
@@
This implies that, in the limit $\kappa\to\infty$, magnetic monopoles cannot move between the boundary and the bulk, i.e. there is a $\cramped{(k_{\mathrm{m}}-1)}$-form symmetry on the boundary (0-form in the Maxwell case) corresponding to conservation of boundary magnetic charge. The action Eq. 115 is precisely the dual of the $\kappa=\infty$ boundary $(k-1)$-form U(1) gauge theory Eq. 100, a 3D XY model in the Maxwell case.
@@
The $\tilde{j}_{\mathrm{m}}$ in the boundary layer must be coupled to $\tilde{b}$ membranes by the magnetic Gauss law, but these have zero tension if they lie entirely within the boundary layer. This means that the Higgs vortices (magnetic field lines) are effectively not present within the boundary layer. This was by construction, since there were no $\mathrm{d}\theta$ terms in the original action which would allow for vortices of the Higgs field within the boundary layer. As a result, in the partition function to leading order the boundary and bulk are effectively decoupled from each other. Given any configuration of monopole worldlines $\tilde{j}_{\mathrm{m}}$, which are closed in the boundary and closed in the bulk, the dominant contribution to the partition function will be for the boundary $\tilde{j}_{\mathrm{m}}$ to be connected to tensionless membranes in the boundary, rather than to be connected to a bulk monopole by a tensionful one. As a result, the monopoles on the boundary can condense at small $\beta$, which is the dual to the boundary symmetry breaking transition we found in the electric formulation. One can then consider turning on small $\kappa^{\prime}$ and doing a strong coupling expansion. This will have the effect of renormalizing the bulk length scale enabling boundary monopoles to extend further into the bulk by extending a magnetic flux tube while preserving the quasi-$(\cramped{D-1})$-dimensional nature of the boundary.
IV.3 Summary and Discussion: Higher-Form Case
@@
In this section, we have generalized our results from Section II on $1$-form Abelian-Higgs models to higher form Abelian-Higgs models in $D$ spacetime dimensions. In particular, we found that $k$-form gauge field coupled to a $(k-1)$-form Higgs field reduces at infinite $\kappa$ to a $(k-1)$-form gauge theory on the boundary whose dynamical degrees of freedom are half-open Wilson branes terminating on the Higgs field as soon as it enters the bulk. This boundary theory exhibits a confinement-deconfinement transition when $\cramped{D>k+2}$ ($\cramped{k_{\text{m}}>0}$), which spontaneously breaks the $(k-1)$-form global matter symmetry at large $\beta$. Table 1 summarizes this pattern of boundary symmetry breaking. In the marginal cases, $\cramped{D=k+2}$ ($\cramped{k_{\text{m}}=0}$), a generalized Mermin-Wagner theorem prevents the symmetry from breaking, except in the case $D=3$ and $k=1$, where we predict a BKT boundary transition. There are also precisely the cases where the same mechanism destabilizes the deconfined phase in the bulk [56, 55]. As in the 1-form Abelian and non-Abelian cases studied numerically in this paper, we expect that this boundary phase transition extends into the phase diagram when bulk fluctuations are restored and will end at a bulk critical point, demarcating a boundary between the Higgs and confining regimes.
@@
The general mechanism for this emergent boundary theory at large $\kappa$ identified by considering higher-form Abelian-Higgs models revolves around the constraint $A=\mathrm{d}\theta$ enforced exactly at infinite $\kappa$. This constraint implies that the Wilson operators which create electric charges attached to electric membranes, $\cramped{\mathrm{exp}[i(\mathrm{d}\hat{\theta}-\hat{A})(M)]}$ for open surfaces $M$, act as the identity. This naïvely indicates that the system is a condensate of electric charge in this limit, i.e. the ground state is a coherent state of the charge annihilation operators. As a consequence, the electric-brane insertion operators $\mathrm{exp}[i\hat{A}(M)]$ for closed surfaces $M$ trivialize in the bulk. In the presence of electric-flux-permeable open boundaries, however, operators inserting electric flux through the boundary are immediately screened as soon as they enter the bulk, terminating on the Higgs field, as shown in Fig. 11, and become charged under the matter symmetry. These operators are the dynamical degrees of freedom at play at the boundary which exhibit the matter symmetry breaking. Presumably when $\kappa$ is reduced the electric flux can penetrate further into the system before being screened by the electric charge condensate, forming a quasi-$(D-1)$-dimensional boundary. It would be interesting to explore whether this mechanism can be extended to higher-form non-Abelian theories described by higher-categorical gauge groups [63, 64, 65, 66, 67].
@@
In the second part of this section, we studied the dualized version of these theories, identifying the boundary degrees of freedom and accounting, in the $1$-form case and $D=4$, for the existence of the dual to the $3$D XY model that we found in terms of direct variables. Magnetic charge moves in the boundary layer, and there is an extra Gauss law in the $\kappa\to\infty$ limit which originates from the constraint that electric charge cannot leave the system. This enforces that magnetic charge cannot leave the boundary into the bulk, and thus the charges on the boundary can condense, leading to the dual phase transition. It would be of interest to study the dual theory in more detail, since it gives a clearer picture of the boundary symmetry breaking in the large-$\kappa$ (small $\kappa^{\prime}$) limit. In particular, it would be worthwhile to pursue Monte Carlo simulations in the dual representation, for which efficient algorithms have been developed [68, 69, 70]. It is also worth re-emphasising the importance of our choice of boundary conditions, which prevented electric charge from leaving the system while allowing electric flux to leave. In the dual theory this led to a flat dual boundary, meaning magnetic charge and flux is always contained in the system and cannot leave. It follows that if one started with flat boundaries, which keep all electric charge and flux inside the system, the dual boundaries would be open, i.e. magnetic charges are kept in the system but magnetic flux can leave. This implies a physical magnetic charge symmetry which can spontaneously break in the confined regime rather than the Higgs regime, with a phase transition on the $\beta=0$ axis instead of the $\kappa^{\prime}=0$ axis.
V Discussion and Conclusion
V.1 Summary and Outlook
@@
In this work we have explored a variety of models of charged Higgs fields coupled to gauge fields in the fundamental representation, under the imposition of boundary conditions which allow electric flux, but not charge, to exit the system. In a closed system, the gauge field does not have a physical global charge symmetry which can spontaneously break, because a charge is always attached to electric flux, which must end inside the system on another charge. With the “electric-flux-permeable” boundary conditions we consider, the charge sectors and global symmetry become physical, as non-zero bulk charge can be compensated by non-zero boundary flux, and can in principle spontaneously break. This work, focusing on the boundary degrees of freedom, complements work exploring the interplay of global and gauge symmetries in the bulk. See, for example, recent work in Refs. [71, 72, 73, 74, 75].
@@
We have considered the Abelian-Higgs model, two types of non-Abelian Higgs models (with fundamental representation and group-valued Higgs fields), and higher-form Abelian-Higgs models. In terms of the inverse gauge coupling $\beta$ and the matter coupling $\kappa$, all of these models share the common feature of a continuity between an electric-charge confining regime at small $\beta,\kappa$, and a Higgs regime at large $\beta,\kappa$, which are two ends of one continuous thermodynamic phase, as proven by Fradkin and Shenker [12]. We have demonstrated through a combination of analytical argument and numerical investigation that, under all but a few marginal cases, there is a boundary phase transition which indicates the spontaneous breaking of the matter symmetry at large $\beta,\kappa$, i.e. in the Higgs regime.
@@
Ref. [19] predicted a boundary phase transition in the case of gauge group $\mathbbm{Z}_{2}$, and verified it using DMRG in $D=2+1$ dimensions. Ref. [20] predicted the boundary transition for a magnetic monopole-free Abelian U(1) Higgs model without numerics. Using lattice gauge theory Monte Carlo simulations, we have explored a wider range of models, both for Abelian and non-Abelian gauge groups. In Section II, we have numerically verified and studied the boundary transition for gauge group U(1) (with monopoles) in $D=3+1$ dimensions, which exhibits a boundary XY transition. In Section III, we extended this to non-Abelian gauge groups, performing numerical simulations for both $\mathrm{SU}(2)$ and $\mathrm{SU}(3)$ gauge theories coupled to fundamental Higgs fields in $D=3+1$, and considered generalizations to $\mathrm{SU}(N)$, $\mathrm{SO}(N)$, and general gauge groups. Lastly, in Section IV we studied higher-form generalizations of Abelian-Higgs models, and demonstrated that the corresponding higher-form matter symmetry can spontaneously break at the boundary. In all of our numerical simulations, the nature of the boundary phase transitions deduced from the numerical data conforms to predictions obtained by studying the $\cramped{\kappa\to\infty}$ limit.
@@
We expect the considerations in this work to extend further to all gauge groups and different Higgs representations. Notably we have not discussed higher-rank gauge theories coupled to scalar matter that are connected to fractonic excitations, but here too one may preserve the Gauss law for tensor fields on the boundary and expect a U(1) global symmetry that, for certain classes of such models, can be broken spontaneously. It would also be of interest to extend these results to discrete non-Abelian gauge groups, and non-Abelian higher-form gauge theories with higher-categorical gauge groups.
Already in the seminal work of Fradkin and Shenker [12] it was understood that in gauge theories with fundamental Higgs matter, under most circumstances thermodynamic quantities exhibit no singularities along paths between Higgs and confined regimes, meaning that they form the same bulk phase of matter. The boundary spontaneous symmetry breaking investigated in this paper does not contradict this result, because it does not define a precise bulk phase boundary between the Higgs and confined regimes. Indeed, as we observed, one may tune the boundary coupling while preserving all the bulk properties to shift the location of the boundary phase transition.
V.2 Higgs=SPT
It would be remiss of us to conclude this paper without a description of some of the work that motivated this study—namely the papers [18, 19, 20] discussing the relationship of Higgs phases to symmetry-protected topological (SPT) phases. We briefly summarize the main findings therein and comment on how our results bear on the generality of this relationship.

### v2

@@ 4.2.1 Open Boundary Duality Transformation
Next, consider the $\cramped{\kappa\to 0}$ ($\cramped{\kappa^{\prime}\to\infty}$) limit. Electric charges are turned off, restoring the electric 1-form symmetry and reducing to a pure U(1) gauge theory. In the dual theory the gauge field $\tilde{H}$ is turned off, $\cramped{\mathrm{d}\tilde{H}=0}$. The theory reduces to a Coulomb gas of magnetic monopole worldlines [31, 66], which has a phase transition separating the deconfined phase (low temperature condensate, 1-form symmetry spontaneously broken) and the confined phase (high temperature gas, 1-form symmetry unbroken).
Lastly, consider the $\cramped{\beta\to 0}$ ($\cramped{\beta^{\prime}\to\infty}$) limit, the strong coupling limit of the original theory. In the electric theory, we can fix unitary gauge to remove $\theta$ and obtain the action $\kappa\cos(A)$ on every link independently, so the system is fully disordered. Equivalently, if we use Eq.˜105, setting $\beta=0$ forces all of the Bessel functions to vanish except when $e=j_{\mathrm{e}}=0$, which reduces the partition function to $\prod_{\ell}I_{0}(\kappa)$. In the dual theory, $\beta^{\prime}\to\infty$, the weak coupling limit, and we have the constraint $\cramped{\tilde{H}=-\mathrm{d}\tilde{A}_{\mathrm{m}}}$, which further implies $\cramped{\mathrm{d}\tilde{H}=0}$. The resulting theory then just has Lagrange multipliers that force the charge loops and membranes to vanish, so the theory trivializes completely.
4.2.2 Dual Boundary Symmetry Breaking
We now consider setting $\kappa^{\prime}=0$ in Eq.˜110 ($\cramped{\kappa\to\infty}$). The variables in play are $\tilde{A}_{\mathrm{m}}$ and $\tilde{j}_{\mathrm{m}}$, defined on all $k_{\mathrm{m}}$-cells of the dual lattice, and $\tilde{H}$ and $\tilde{b}$ on all ($k_{\mathrm{m}}+1$)-cells. In the bulk, the $\kappa^{\prime}=0$ means that the kinetic term for $\tilde{H}$ drops out and so the fluxes of $\tilde{H}$ (corresponding to electric charges of the original action) are completely unconstrained. In other words, the bulk $\tilde{H}$ is in the strong coupling limit. We may integrate out $\tilde{H}$, to obtain
@@ 4.2.2 Dual Boundary Symmetry Breaking
Now consider an open boundary. Recall that the boundary of the dual lattice is “flat”, as shown in Fig.˜11, i.e. it has no cells extending into the vacuum. This means that no magnetic charge or magnetic flux can exit the system. More concretely, the constraint $\cramped{j_{\mathrm{e}}|_{\partial X}=0}$, in Eq.˜104, in the electric variables is reflected in the dual constraint $\cramped{\mathrm{d}\tilde{H}|_{\partial\tilde{X}}=0}$, which enforces that $\tilde{H}$ is pure gauge in the dual boundary layer. This means that the boundary action has no $\kappa$ dependence (it effectively has $\kappa=0$). We resolve the boundary constraint as $\tilde{H}=\mathrm{d}\tilde{f}$, so that
@@
We then define a composite field $\tilde{\chi}=\tilde{A}_{\mathrm{m}}+\tilde{f}$, which is gauge invariant under the gauge transformations $\tilde{A}_{\mathrm{m}}\to\tilde{A}_{\mathrm{m}}+\tilde{\lambda}$ and $f\to\tilde{f}-\tilde{\lambda}$. This allows us to rewrite the boundary action as
@@
To proceed from here we need to be careful about how the $\tilde{j}_{\mathrm{m}}$ and $\tilde{b}$ can move between the bulk and boundary layers. We do so in the $\cramped{\kappa\to\infty}$ limit, by combining this with the bulk action Eq.˜113. We can then integrate out $\tilde{A}_{\mathrm{m}}$ to generate the magnetic Gauss law, which is enforced on every link. This leaves us with the total action
$S_{\text{dual}}\xrightarrow{\kappa\to\infty}-\frac{1}{2\beta^{\prime}}\langle\tilde{b},\tilde{b}\rangle_{\tilde{X}}-\frac{\beta^{\prime}}{2}(\mathrm{d}\tilde{\chi})_{\partial\tilde{X}}^{2}-i\langle\tilde{\chi},\tilde{j}_{\mathrm{m}}|_{\partial\tilde{X}}\rangle$. (117)
@@
One should be concerned here as to how the $\tilde{j}_{\mathrm{m}}$ from the bulk (edges of $\tilde{b}$) couple to the boundary. The key to understand what happens here is that (i) the $\tilde{b}$ membranes can lie in the boundary layer where they cost zero action, and (ii) for this action to be gauge-invariant under shifts of $\tilde{\chi}$, we must have an additional boundary Gauss law,
$\mathrm{d}^{\dagger}\tilde{j}_{\mathrm{m}}|_{\partial\tilde{X}}=0$ (118)
@@
This implies that, in the limit $\cramped{\kappa\to\infty}$, magnetic monopoles cannot move between the boundary and the bulk, i.e. there is a $\cramped{(k_{\mathrm{m}}-1)}$-form symmetry on the boundary (0-form in the Maxwell case) corresponding to conservation of boundary magnetic charge. The action Eq.˜117 is precisely the dual of the $\cramped{\kappa=\infty}$ boundary $\cramped{(k-1)}$-form U(1) gauge theory Eq.˜102, a 3D XY model in the Maxwell case.
@@
The $\tilde{j}_{\mathrm{m}}$ in the boundary layer must be coupled to $\tilde{b}$ membranes by the magnetic Gauss law, but these have zero tension if they lie entirely within the boundary layer. This means that the Higgs vortices (magnetic field lines) are effectively not present within the boundary layer. This was by construction, since there were no $\mathrm{d}\theta$ terms in the original action which would allow for vortices of the Higgs field within the boundary layer. As a result, in the partition function to leading order the boundary and bulk are effectively decoupled from each other. Given any configuration of monopole worldlines $\tilde{j}_{\mathrm{m}}$, which are closed in the boundary and closed in the bulk, the dominant contribution to the partition function will be for the boundary $\tilde{j}_{\mathrm{m}}$ to be connected to tensionless membranes in the boundary, rather than to be connected to a bulk monopole by a tensionful one. As a result, the monopoles on the boundary can condense at small $\beta$, which is the dual to the boundary symmetry breaking transition we found in the electric formulation.
One can then consider turning on small $\kappa^{\prime}$ and doing a strong coupling expansion. This will have the effect of renormalizing the bulk length scale enabling boundary monopoles to extend further into the bulk by extending a magnetic flux tube while preserving the quasi-$(\cramped{D-1})$-dimensional nature of the boundary.
4.3 Summary and Discussion: Higher-Form Case
@@
In this section, we have generalized our results from Section˜2 on $1$-form Abelian-Higgs models to higher form Abelian-Higgs models in $D$ spacetime dimensions. In particular, we found that $k$-form gauge field coupled to a (k1)(k-1)-form Higgs field reduces at infinite $\kappa$ to a (k1)(k-1)-form gauge theory on the boundary whose dynamical degrees of freedom are half-open Wilson branes terminating on the Higgs field as soon as they enter the bulk. This boundary theory exhibits a confinement-deconfinement transition when $\cramped{D>k+2}$ ($\cramped{k_{\text{m}}>0}$), which spontaneously breaks the (k1)(k-1)-form global matter symmetry at large $\beta$. Table 1 summarizes this pattern of boundary symmetry breaking. In the marginal cases, $\cramped{D=k+2}$ ($\cramped{k_{\text{m}}=0}$), a generalized Mermin-Wagner theorem prevents the symmetry from breaking, except in the case $D=3$ and $k=1$, where we predict a BKT boundary transition. There are also precisely the cases where the same mechanism destabilizes the deconfined phase in the bulk [64, 63]. As in the 1-form Abelian and non-Abelian cases studied numerically in this paper, we expect that this boundary phase transition extends into the phase diagram when bulk fluctuations are restored and will end at a bulk critical point, demarcating a boundary between the Higgs and confining regimes.
@@
The general mechanism for this emergent boundary theory at large $\kappa$ identified by considering higher-form Abelian-Higgs models revolves around the constraint $A=\mathrm{d}\theta$ enforced exactly at infinite $\kappa$. This constraint implies that the Wilson operators which create electric charges attached to electric membranes, $\cramped{\mathrm{exp}[i(\mathrm{d}\hat{\theta}-\hat{A})(M)]}$ for open surfaces $M$, act as the identity. This naïvely indicates that the system is a condensate of electric charge in this limit, i.e. the ground state is a coherent state of the charge annihilation operators. As a consequence, the electric-brane insertion operators $\mathrm{exp}[i\hat{A}(M)]$ for closed surfaces $M$ trivialize in the bulk. In the presence of electric-flux-permeable open boundaries, however, operators inserting electric flux through the boundary are immediately screened as soon as they enter the bulk, terminating on the Higgs field, as shown in Fig.˜12, and become charged under the matter symmetry. These operators are the dynamical degrees of freedom at play at the boundary which exhibit the matter symmetry breaking. Presumably when $\kappa$ is reduced the electric flux can penetrate further into the system before being screened by the electric charge condensate, forming a quasi-(D1)(D-1)-dimensional boundary. It would be interesting to explore whether this mechanism can be extended to higher-form non-Abelian theories described by higher-categorical gauge groups [70, 71, 72, 73, 74].
@@
In the second part of this section, we studied the dualized version of these theories, identifying the boundary degrees of freedom and accounting, in the $1$-form case and $D=4$, for the existence of the dual to the 33D XY model that we found in terms of direct variables. Magnetic charge moves in the boundary layer, and there is an extra Gauss law in the $\cramped{\kappa\to\infty}$ limit which originates from the constraint that electric charge cannot leave the system. This enforces that magnetic charge cannot leave the boundary into the bulk, and thus the charges on the boundary can condense, leading to the dual phase transition. It would be of interest to study the dual theory in more detail, since it gives a clearer picture of the boundary symmetry breaking in the large-$\kappa$ (small $\kappa^{\prime}$) limit. In particular, it would be worthwhile to pursue Monte Carlo simulations in the dual representation, for which efficient algorithms have been developed [75, 76, 77]. It is also worth re-emphasising the importance of our choice of boundary conditions, which prevented electric charge from leaving the system while allowing electric flux to leave. In the dual theory this led to a flat dual boundary, meaning magnetic charge and flux is always contained in the system and cannot leave. It follows that if one started with flat boundaries, which keep all electric charge and flux inside the system, the dual boundaries would be open, i.e. magnetic charges are kept in the system but magnetic flux can leave. This implies a physical magnetic charge symmetry which can spontaneously break in the confined regime rather than the Higgs regime, with a phase transition on the $\beta=0$ axis instead of the $\kappa^{\prime}=0$ axis. [^11] This was also discussed in the $\mathbbm{Z}_{2}$ 1-form case in [22].
5 Discussion and Conclusion
5.1 Summary and Outlook
@@
In this work we have explored a variety of models of charged Higgs fields coupled to gauge fields in the fundamental representation, under the imposition of boundary conditions which allow electric flux, but not charge, to exit the system. In a closed system, the gauge field does not have a physical global charge symmetry which can spontaneously break, because a charge is always attached to electric flux, which must end inside the system on another charge. With the “electric-flux-permeable” boundary conditions we consider, the charge sectors and global symmetry become physical, as non-zero bulk charge can be compensated by non-zero boundary flux, and can in principle spontaneously break. This work, focusing on the boundary degrees of freedom, complements work exploring the interplay of global and gauge symmetries in the bulk. See for example recent work in Refs. [78, 79, 80, 81, 82]. More broadly, the boundary perspective offers novel insight into the gauge-invariant description of Higgs phases and how they might be distinguished from confined phases, issues with a long history [15, 14, 13, 25, 26] which continue to generate interest to the present day [17, 83, 37, 19, 20, 21, 22, 23, 84, 85, 18].
@@
We have considered the Abelian-Higgs model, two types of non-Abelian Higgs models (with fundamental representation and group-valued Higgs fields), and higher-form Abelian-Higgs models. In terms of the inverse gauge coupling $\beta$ and the matter coupling $\kappa$, all of these models share the common feature of a continuity between an electric-charge confining regime at small $\beta,\kappa$, and a Higgs regime at large $\beta,\kappa$, which are two ends of one continuous thermodynamic phase, as proven by Fradkin and Shenker [14]. We have demonstrated through a combination of analytical argument and numerical investigation that, under all but a few marginal cases, there is a boundary phase transition which indicates the spontaneous breaking of the matter symmetry at large $\beta,\kappa$, i.e. in the Higgs regime, as illustrated in Fig.˜1. Already in the seminal work of Fradkin and Shenker [14] it was understood that in gauge theories with fundamental Higgs matter, under most circumstances thermodynamic quantities exhibit no singularities along paths between Higgs and confined regimes, meaning that they form the same bulk phase of matter. The boundary spontaneous symmetry breaking investigated in this paper does not contradict this result, because it does not define a precise bulk phase boundary between the Higgs and confined regimes. Indeed, as we observed, one may tune the boundary coupling while preserving all the bulk properties to shift the location of the boundary phase transition.
@@
Ref. [22] predicted a boundary phase transition in the case of gauge group $\mathbbm{Z}_{2}$, and verified it using DMRG in $D=2+1$ dimensions. Ref. [23] predicted the boundary transition for a magnetic monopole-free Abelian U(1) Higgs model without numerics. Using lattice gauge theory Monte Carlo simulations, we have explored a wider range of models, both for Abelian and non-Abelian gauge groups. In Section˜2, we have numerically verified and studied the boundary transition for gauge group U(1) (with monopoles) in $D=3+1$ dimensions, which exhibits a boundary XY transition. In Section˜3, we extended this to non-Abelian gauge groups, performing numerical simulations for both $\mathrm{SU}(2)$ and $\mathrm{SU}(3)$ gauge theories coupled to fundamental Higgs fields in $D=3+1$, and considered generalizations to $\mathrm{SU}(N)$, $\mathrm{SO}(N)$, and general gauge groups. Lastly, in Section˜4 we studied higher-form generalizations of Abelian-Higgs models, and demonstrated that the corresponding higher-form matter symmetry can spontaneously break at the boundary. In all of our numerical simulations, the nature of the boundary phase transitions deduced from the numerical data conforms to predictions obtained by studying the $\cramped{\kappa\to\infty}$ limit.
@@
We expect the considerations in this work to extend further to all gauge groups and different Higgs representations [52]. Notably we have not discussed higher-rank gauge theories coupled to scalar matter that are connected to fractonic excitations [86, 87, 88, 89, 90, 91], but here too one may preserve the Gauss law for tensor fields on the boundary and expect a U(1) global symmetry that, for certain classes of such models, can be broken spontaneously. It would also be of interest to extend these results to discrete non-Abelian gauge groups, and non-Abelian higher-form gauge theories with higher-categorical gauge groups [71, 72, 73, 74].
Before continuing, we comment briefly on some interesting connections between our work and two other active areas of research: boundary criticality and asymptotic symmetries. “Traditional” boundary criticality has a long history [92, 93], and has recently received some new insights [94, 95]. The central idea is that the boundary of a system may undergo a phase transition in a novel “extraordinary” universality class when coupled to a critical bulk, by tuning the boundary coupling relative to the bulk. In this work we have only considered the behavior of the boundary when the bulk is gapped, but it would be interesting to explore the boundary criticality in more detail in the vicinity of the bulk critical point where the transition appears to end for $\alpha\geq 1$ (cf. Figs.˜4 and 8), especially at the lower critical dimension [94]. Prior boundary criticality studies have focused on criticality from the spontaneously breaking of 0-form symmetries, but our work raises the possibility of novel boundary physics arising from dynamical gauge fields and higher-form symmetries. On the one hand, such exploration may yield novel boundary universality classes, and on the other it may yield some new insights into the behavior of these enigmatic bulk critical points [96, 97].
Another active area with some overlap with this work is the exploration of asymptotic symmetries in gauge theories and gravity on the boundary of compactified spacetime [98, 99, 100, 101, 102, 103, 104, 105, 106, 107]. For example, in pure electromagnetism in Minkowski spacetime at null infinity it is well known that there is an infinite dimensional symmetry generated by large gauge transformations with conserved charges living on the boundary [100]. Indeed, Ref. [103] has already made some direct connections between asymptotic (or “long-range gauge”) symmetries and the boundary symmetries studied in this work. It would be interesting to deepen the connections between Higgs boundary criticality on the lattice and asymptotic symmetries in the continuum.
5.2 Higgs = SPT